# Lightweight Gradient-Aware Upscaling of 3D Gaussian Splatting Images

### v1

@@ Abstract
We introduce an image upscaling technique tailored for 3D Gaussian Splatting (3DGS) on lightweight GPUs. Compared to 3DGS, it achieves significantly higher rendering speeds and reduces artifacts commonly observed in 3DGS reconstructions. Our technique upscales low-resolution 3DGS renderings with a marginal increase in cost by directly leveraging the analytical image gradients of Gaussians for gradient-based bicubic spline interpolation. The technique is agnostic to the specific 3DGS implementation, achieving novel view synthesis at rates 3×–4× higher than the baseline implementation. Through extensive experiments on multiple datasets, we showcase the performance improvements and high reconstruction fidelity attainable with gradient-aware upscaling of 3DGS images. We further demonstrate the integration of gradient-aware upscaling into the gradient-based optimization of a 3DGS model and analyze its effects on reconstruction quality and performance.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) [14] has rapidly become one of the most popular methods for novel view synthesis, offering an efficient representation that closely fits recorded scenes and can be rendered efficiently using GPU rasterization. Despite continuous improvements in reconstruction [33, 27] and rendering performance [23], powerful GPUs are still required to efficiently optimize and render 3DGS models from high-resolution images. For instance, at 4K resolution, when capturing complex scenes such as the bicycle scene from the Mip-NeRF 360 dataset [3], even high-end GPUs like NVIDIA’s A5000 struggle to maintain a stable 30 FPS.
This limitation becomes even more critical in VR applications, where two high-resolution images must be rendered simultaneously at a high frame rate to maintain a smooth and immersive experience. In scenarios where 3DGS models are accessed via mobile devices with high-resolution screens but lightweight GPUs [22], rendering performance often drops to an unacceptably low level.
A common strategy to enhance image details and make images suitable for higher-resolution displays without increasing the rasterization load is image upscaling. Image upscaling methods such as Bicubic and Lanczos interpolation are widely used [1] but often introduce noticeable artifacts, reducing visual quality. Deep learning (DL)-based upscalers offer superior quality by preserving and enhancing image details, but they also have significant drawbacks. Approaches like NinaSR [9] or SRGS [7] are not fast enough for real-time applications, while solutions such as NVIDIA DLSS [24] require specialized hardware and rely on the use of additional image information, such as accurate depth buffers and motion vectors, which are unavailable for 3DGS. All DL-based approaches may hallucinate incorrect details (e.g., adding noise or sharpening areas incorrectly), which is especially problematic in VR, where non-consistent upscaling for both eyes creates an unpleasant experience.
We propose an image upscaling technique specifically tailored for 3DGS to address these limitations. It outperforms classical upscaling techniques regarding quality, such as Lanczos interpolation, while introducing low computational overhead. We employ a spline-based bicubic interpolation formulation with the Gaussians’ analytical image gradients to achieve this.
Compared to 3DGS, our technique achieves significantly improved frame rates at high fidelity (see Fig. 1). The quality-improved variant of 3DGS, Mip-Splatting [33], achieves similar quality but has limited potential for real-time applications as it further increases the rasterization workload (see supplemental material).
@@ 5 Image Gradient Based Upscaling
In the 3DGS forward rendering pass, the analytical gradients are available at every pixel of the used pixel raster. These gradients are then used as outlined to perform spline-based image interpolation. The 3DGS rendering process, including image upscaling, is illustrated in Fig. 4.
@@
Since both the image gradient calculation and spline-based interpolation are fully differentiable with respect to the scene parameters, image upscaling using analytical gradients can be seamlessly integrated into the optimization pipeline for reconstructing the 3DGS model. Fig. 4 illustrates the adapted reconstruction pipeline. For bicubic interpolation, the pipeline remains identical, except that gradients are computed using finite differences during the interpolation stage.
@@ C Comparison to Mip Splatting
Since our primary goal is efficient rendering, we excluded Mip-Splatting from the main paper. While Mip Splatting improves image consistency at high resolutions, its computational overhead makes it less suitable for real-time applications on low-end devices. Nevertheless, our upscaling method can be applied to scenes reconstructed with Mip Splatting, thereby speeding up the rendering process.

### v2

@@ Abstract
We introduce an image upscaling technique tailored for 3D Gaussian Splatting (3DGS) on lightweight GPUs. Compared to 3DGS, it achieves significantly higher rendering speeds and reduces commonly observed reconstruction artifacts. Our technique upscales low-resolution 3DGS renderings with a marginal increase in cost by directly leveraging the analytical image gradients of Gaussians for gradient-based bicubic spline interpolation. The technique is agnostic to the specific 3DGS implementation, achieving novel view synthesis at rates 3×–4× higher than the baseline implementation. Through extensive experiments on multiple datasets, we showcase the performance improvements and high reconstruction fidelity attainable with gradient-aware upscaling of 3DGS images. We further demonstrate the integration of gradient-aware upscaling into the gradient-based optimization of a 3DGS model and analyze its effects on reconstruction quality and performance.
@@ 1 Introduction
Despite continuous improvements of 3D Gaussian Splatting (3DGS) [14] in reconstruction [34, 28] and rendering performance [24], even high-end GPUs struggle to maintain a stable 30 FPS when rendering to high-resolution displays. This limitation is even more critical in VR applications, where two high-resolution images must be rendered simultaneously to maintain a smooth and immersive experience. When 3DGS models are accessed via mobile devices with high-resolution screens but lightweight GPUs [23], rendering performance often drops to an unacceptably low level.
A common strategy to render to higher-resolution displays without increasing the rasterization load is image upscaling. Image upscaling methods such as Bicubic and Lanczos interpolation are widely used [1] but often reduce visual quality by introducing noticeable artifacts. Deep learning (DL)-based upscalers offer superior quality by enhancing image details, but they also have significant drawbacks. High-quality approaches like NinaSR [9], SRGS [7] or SwinIR [18] are not fast enough for real-time applications, while solutions such as NVIDIA DLSS [25] require specialized hardware and rely on the use of additional image information, such as accurate depth buffers and motion vectors, which are unavailable for 3DGS. Furthermore, when using DL-based upscalers in VR, users often report issues due to temporal inconsistencies and feature popping. Although such upscalers produce sharp images, the results often appear unnatural due to hallucinated artifacts, and PSNR scores tend to decrease (see Fig. 2).
We propose a computationally lightweight image upscaling technique specifically tailored for 3DGS to address these limitations. Innovation comes from uniquely leveraging Gaussian primitives’ gradients, introducing a new way of thinking about upscaling: not using fixed kernels as in most classical methods, not requiring heavy pre-training and high upscaling times as modern neural approaches, and delivering faster upscaling with superior image quality than highly optimized GPU-based upscalers.
Compared to 3DGS, we achieve significantly improved frame rates at high fidelity. The quality-improved variant of 3DGS, Mip-Splatting [34], shows limited quality reconstruction of regions sparsely viewed in the test set and further increases the rasterization workload (see supplemental material).
@@ 5 Image Gradient Based Upscaling
In the 3DGS forward rendering pass, the analytical gradients are available at every pixel of the used pixel raster. These gradients are then used as outlined to perform spline-based image interpolation. The fully differentiable 3DGS rendering process , including image upscaling, is illustrated in Fig. 5.
@@
Since both the image gradient calculation and spline-based interpolation are fully differentiable with respect to the scene parameters, image upscaling using analytical gradients can be seamlessly integrated into the optimization pipeline for reconstructing the 3DGS model. For bicubic interpolation, the pipeline remains identical, except that gradients are computed using finite differences during the interpolation stage.
@@ C Comparison to Mip Splatting
While Mip Splatting improves image consistency at high resolutions, its computational overhead makes it less suitable for real-time applications on low-end devices. Nevertheless, our upscaling method can be applied to scenes reconstructed with Mip Splatting, thereby speeding up the rendering process.